# BugRepro: Enhancing Android Bug Reproduction with Domain-Specific Knowledge Integration

### v1

@@ Abstract
Mobile application development is a fast-paced process where maintaining high-quality user experiences is crucial. Current bug reproduction methods predominantly depend on precise feature descriptions in bug reports. However, the growing complexity and dynamism of modern software systems pose significant challenges to this crucial quality assurance process, as ambiguous or incomplete steps-to-reproduce (S2Rs) in reports frequently impede effective debugging and maintenance. To address these challenges, we propose BugRepro, a novel technique that integrates domain-specific knowledge to enhance the accuracy and efficiency of bug reproduction. BugRepro adopts a Retrieval-Augmented Generation (RAG) approach. It retrieves similar bug reports along with their corresponding S2R entities from an example-rich RAG document. This document serves as a valuable reference for improving the accuracy of S2R entity extraction. In addition, BugRepro incorporates app-specific knowledge. It explores the app’s graphical user interface (GUI) and extracts UI transition graphs. These graphs are used to guide large language models (LLMs) in their exploration process when they encounter bottlenecks. Our experiments demonstrate the effectiveness of BugRepro. Our method significantly outperforms two state-of-the-art methods. For S2R entity extraction accuracy, it achieves improvements of 8.85% and 28.89%. For bug reproduction success rate, the improvements reach 74.55% and 152.63%. In reproduction efficiency, the gains are 0.72% and 76.68%.
@@ 1 Introduction
Mobile applications have become an essential part of daily life, with over 4 million apps available on Google Play and the Apple App Store as of 14 December 2024 [1]. As the market for mobile apps continues to grow, ensuring a seamless user experience has become a critical priority for developers striving to remain competitive. A key aspect of maintaining app quality is addressing the issues reported by users, typically through bug reports. These reports often include steps to reproduce (S2R), enabling developers to replicate the reported problems, identify their root causes, and implement fixes.
@@
Existing methods still heavily rely on the original content of bug reports. Specifically, when the descriptions are ambiguous or difficult to comprehend, current approaches fail to extract accurate and useful information from these reports. Moreover, modern applications exhibit inherent complexity with multiple pages and diverse functionalities, making it challenging for existing methods to map the relevant information in bug reports to the corresponding UI elements that need to be manipulated.
@@
To overcome these constraints, we propose BugRepro, a novel technique that integrates bug-report-specific knowledge and app-specific knowledge to enhance the bug reproduction process. Specifically, BugRepro first employs a Retrieval-Augmented Generation (RAG) [33] approach to retrieve similar bug reports and their corresponding S2R entities from an example-rich RAG document. The retrieved knowledge serves as a reference, enabling LLMs to extract S2R entities more accurately for the current bug report. Next, BugRepro incorporates app-specific knowledge by exploring the app’s GUI to improve the LLM-guided bug reproduction process further. When LLMs encounter challenges, such as difficulty in determining which UI component to interact with or missing steps in the reproduction process, BugRepro extracts UI transition graphs(UTG) that model the changes and interactions of the app state on different screens. These graphs are then fed into the LLM to guide its decision-making, ensuring the reproduction follows the correct sequence of actions and app state transitions.
We conduct extensive experiments on 151 real-world bug reports to evaluate the performance of BugRepro. First, we evaluate the accuracy of S2R extraction and compare BugRepro with two state-of-the-art techniques (i.e., RecDroid and AdbGPT ). The experimental results demonstrate that BugRepro outperforms these approaches, achieving improvements of 7.57% and 28.89%. in terms of the extraction accuracy, respectively. Then, we evaluate the effectiveness of bug replay. The experimental results show that RecDroid and AdbGPT successfully reproduced 38 and 55 bugs, whereas BugRepro can reproduce 96 bugs, achieving an improvement of 152.63% and 74.55%, respectively. Furthermore, BugRepro reduces the time required for bug reproduction, with an average time of 124.7 seconds, compared to 534.9 seconds for RecDroid and 125.6 seconds for AdbGPT. Finally, ablation studies confirm that each component of BugRepro contributes to the overall performance.
@@
  • BugRepro effectively addresses three key challenges in Android bug reproduction: the lack of domain-specific knowledge, the complexity of modern applications with multi-page structures and diverse functionalities, and the instability of LLM-generated output formats in the context of code interaction.
@@
2 Motivation and Background
@@ 2.1 Motivation
Table 1 provides examples of real-world bug reports, and Table 2 presents a comparison of the corresponding S2R entity extraction results.
Table 1: An example of Bug Reports
@@
We observe that prior approaches tend to heavily rely on the presence of explicit action verbs(e.g., “Tap”, “Click”). Consequently, they often fail to extract S2R entities when action verbs are implicit or absent in the sentence. Moreover, these methods frequently fail to extract entities when sentences contain multiple components (e.g., “Secret field” or “other required fields”) or actions (e.g., “enter” or “fill”) of the same type, resulting in incomplete or ambiguous extraction results.
During the replay stage, prior methods [57, 16] share the same shortcoming—neither nor deal with scenarios that involve multi-page flows and complex functionalities. According to the algorithms used in prior approaches, during the exploration process, if a bug report skips the description of a component on the current screen and directly mentions a component on the next screen, the system tends to perform a “back” operation to return to the previous screen. For example, a bug report says “Select About”, but on the current screen, there is no component labeled “About”. We need to select any component from this screen first, and then we can see the “About” component. When prior methods encounter this situation, they simply choose the “back” action, which never leads to the correct screen. This often leads to incorrect looping behavior, ultimately causing the failure of bug reproduction.
2.2 Large Language Models
2.3 Retrieval-Augmented Generation
2.4 Bug Reports and App UI Elements
@@ 3 Approach
Our method is structured into two distinct phases and comprises three core components.The first phase is RAG-enhanced S2R entity extraction, we employ RAG to enhance the accuracy of S2R entity extraction 3.1.2. When a new bug report is submitted, it is first segmented into individual sentences. For each sentence, the system automatically retrieves semantically similar sentences and their corresponding labels (which include all associated S2R entities) from the RAG database—an example is shown in Table 3. Then, the original bug report along with the retrieved labeled examples is fed into LLM, which is guided by the examples to analyze and extract the S2R entities corresponding to the new bug report. After extracting the S2R entities, we leverage LLMs to interact with UI components based on these entities and corresponding bug reports to replay the bug. At the same time, based on existing work and our observations, we found that even after applying RAG, challenges remain in extracting components, which may lead to difficulties in correctly matching and operating the corresponding components during the replay process. Therefore, we need to provide the LLM with more UI-related knowledge. the second phase is Exploration-based Replay, which aims to provide app-related knowledge to LLMs, enabling them to gain insights into the app, understand its usage, and make informed decisions. It consists of two primary components: UI exploration based on functionality and LLM guided replay, which continuously interact with each other throughout the process. The functionality-aware UI exploration component based on functionality is responsible for exploring the UI, and all information will be transmitted to LLM. Note that since UI screens are visual and can not be processed by LLMs directly, we encode the UI screens into text format for LLMs processing following existing work [16]. The LLM guided replay component receives the encoded UI information from the functionality-aware UI exploration module, synthesizes it with inputs from the RAG-enhanced S2R entity extraction module and the original bug report, and then generates guidance for the next interaction step based on this comprehensive context. After executing each step, Bugrepro checks whether the bug reproduction has succeeded and provides feedback to the LLM. If reproduction fails, it triggers the functionality-aware UI exploration process, allowing the acquisition of additional domain-specific knowledge to enhance the LLM’s decision-making. During the LLM’s decision-making process, when the LLM faces difficulties determining the next step (§ 3.2.1), BugRepro augments its decision-making by integrating app-specific knowledge, which encompasses both UI exploration and the synthesis of app-related information. This process involves analyzing the app’s UI elements and interactions to construct the Synthesized Functionality Table (Table 5), which consists of three components: <Synthesized functionality, UI states, UI elements>. Once equipped with more comprehensive domain-specific knowledge, the LLM is able to make more accurate and effective decisions in response to the current situation.
@@ 3.1.2 RAG-enhanced S2R entity extraction
After defining the available actions and their corresponding primitives, BugRepro incorporates the retrieved examples into the prompt to further assist the LLM. Specifically, BugRepro includes each retrieved sentence along with its corresponding annotated S2R entity, formatted as follows: Here are some examples for S2R entity extraction. The sentence is “Click on the add to option and select playlist.”, the extracted S2R entities are: 1. [Tap] [add to option], 2. [Tap] [playlist].
@@ 3.2.1 UTG-Driven Functional Exploration
UI exploration is triggered when the LLM encounters uncertainty regarding which UI component to interact with. Therefore, BugRepro begin the exploration by analyzing the UTG starting from the page where the difficulty arises, denoted as $p$. The UTG, created by the UI Automator, provides essential information about the app, including the relationships between UI states and the presence of various UI elements on each screen. By examining the functionalities of these elements, we can uncover the tasks that can be performed within the app and identify the UI components necessary to execute them. Therefore, BugRepro parses the UI states and elements in the UTG, querying LLMs to extract their respective functions.
@@ 3.2.2 Exploration-enhanced Prompting
However, when the LLM struggles to make the correct decision—such as when it cannot determine which UI component to interact with, or when exploration based on the LLM’s guidance fails to reproduce the bug and the LLM does not update its decision after revisiting the page—we provide additional app-related knowledge obtained from the previous exploration to assist the LLM. Specifically, given a problematic page, we supply the LLM with the functionality of all UI components on that page and the functionalities of the new UI screens they can navigate to. This additional information enables the LLM to decide which UI component to interact with and what action to take. At this stage, the prompt includes the following components: (1) the bug report text, (2) the extracted S2R entity, (3) the encoded text of the current UI screen, and (4) the app-related knowledge, which includes the synthesized functionalities (§ 3.2.1) of each UI component within current UI page, and the UI function table, which includes a mapping of functionalities associated with UI pages accessible through specific components on the current page.
@@ 4.2 Implementation and Environment
We implement BugRepro in Python. We leverage all-MiniLM- L12-v2 model as the retrieval model used in the RAG component, which is based on the Dense Passage Retrieval (DPR) [30] architecture and uses a bi-encoder to encode the queries and documents into dense vectors. We utilize the DeepSeek-V3 model as the underlying LLM in BugRepro, accessed via its API service [13]. To address potential verbosity in the LLM’s output—such as repeated questions or chain-of-thought reasoning—we adopt a filtering mechanism that removes non-JSON content, retaining only valid JSON segments enclosed in square brackets. For interacting with UI widgets on the device, we leverage UI Automator2 [39] as the execution engine. We set the UTG exploration depth to 1 and the time budget to 5 minutes to balance effectiveness and efficiency. All experiments are performed on a workstation running Windows 11 Pro 64-bit, equipped with a 13th Gen Intel(R) Core(TM) i7-13700KF CPU (24 cores, base frequency  3.4GHz) and an NVIDIA GeForce RTX 4090 GPU.
@@ 5.1 RQ1: How accurate is BugRepro in extracting S2R entities?
From the table, BugRepro outperforms both ReCDroid and AdbGPT across all dimensions, with accuracy improvements ranging from 8.85% to 14.47% compared to AdbGPT. For instance, BugRepro records 69.49% accuracy in extracting action types, compared to 61.92% by AdbGPT and 45.50% by ReCDroid. Similarly, in identifying target components, BugRepro achieves 43.24%, outperforming AdbGPT (28.77%) and ReCDroid (4.98%). This trend persists across input values and scroll directions, underscoring BugRepro’s consistent advantage.
@@
RQ1 summary: Integrating RAG into our approach clearly demonstrates that BugRepro outperforms baseline methods in S2R extraction, achieving accuracy improvements of 8.85% to 14.47% over AdbGPT (state-of-the-art baselines) across all dimensions. This fully confirms that RAG not only provides contextualized examples, enabling BugRepro to effectively handle ambiguous or complex bug reports, but also enhances contextual understanding, precision in multi-step scenarios, and robustness to linguistic variations.
@@ 5.4 RQ4: How does the robustness of each method vary across different language models?
To answer RQ4, we evaluate BugRepro’s performance using both GPT-4 and our default DeepSeek model to assess its generalization capability. Table 10 presents the results, demonstrating consistent S2R entity extraction performance across both models, with DeepSeek showing slight advantages in Action (72.32% vs. 69.60%) and Component (37.60% vs. 33.76%) extraction. For replay performance, DeepSeek achieves a higher NSR of 96 compared to GPT-4’s 86, though with a longer average reproduction time (124.7s vs. 83.0s).
@@
6 Discussion
The primary construct threats stem from the selection of evaluation metrics and the inherent randomness in LLM outputs. LLMs frequently produce unstructured responses, which impedes their integration with structured code and automation pipelines. To address these challenges, we initially adopt approaches from prior work [17], which attempt to enforce output formatting through prompt examples (e.g., “1. [Tap] [‘bookmark’]”). However, such efforts have not fully standardized the responses. As illustrated in Table 11, LLM outputs often display inconsistencies in format, structure, and verbosity, thereby complicating automation and diminishing accuracy.
Our approach employs LangChain to enforce output constraints and ensure standardized formatting. The standardized outputs are shown in Table 6. The refinement of the output format further boosts the performance of BugRepro, enforces structured action output an ensure consistency and automation readiness.

### v2

@@ Abstract
Mobile application development is a fast-paced process where maintaining high-quality user experiences is crucial. Bug reproduction, a key aspect of maintaining app quality, often faces significant challenges. Specifically, when descriptions in bug reports are ambiguous or difficult to comprehend, current approaches fail to extract accurate information. Moreover, modern applications exhibit inherent complexity with multiple pages and diverse functionalities, making it challenging for existing methods to map the relevant information in bug reports to the corresponding UI elements that need to be manipulated. To address these challenges, we propose BugRepro, a novel technique that integrates domain-specific knowledge to enhance the accuracy and efficiency of bug reproduction. BugRepro adopts a Retrieval-Augmented Generation (RAG) approach. It retrieves similar bug reports along with their corresponding steps to reproduce (S2R) entities from an example-rich RAG document. In addition, BugRepro explores the graphical user interface (GUI) of the app and extracts transition graphs from the user interface to incorporate app-specific knowledge to guide large language models (LLMs) in their exploration process. Our experiments demonstrate that BugRepro significantly outperforms two state-of-the-art methods (ReCDroid and AdbGPT). For S2R entity extraction accuracy, it achieves a 7.57 to 28.89 percentage point increase over prior methods. For the bug reproduction success rate, the improvement reaches 74.55% and 152.63%. In reproduction efficiency, the gains are 0.72% and 76.68%.
@@ 1 Introduction
Mobile applications have become an essential part of daily life, with over 4 million apps available on Google Play and the Apple App Store as of 14 December 2024 [1]. As the market for mobile apps continues to grow, ensuring a seamless user experience has become a critical priority for developers striving to remain competitive. A key aspect of maintaining app quality is addressing the issues reported by users, typically through bug reports. These reports often include S2R entities, enabling developers to replicate the reported problems, identify their root causes, and implement fixes.
@@
Although prior approaches have made significant contributions to improving bug reproduction efficiency, critical limitations remain unresolved. We crystallize these challenges as follows:
@@
To overcome these constraints, we propose BugRepro, a novel technique that integrates bug-report-specific knowledge and app-specific knowledge to enhance the bug reproduction process. Specifically, to overcome the first challenge, BugRepro first employs a Retrieval-Augmented Generation (RAG) [33] approach to retrieve similar bug reports and their corresponding S2R entities from an example-rich RAG document. The retrieved knowledge serves as a reference, addressing the limitation of existing LLMs—their lack of domain understanding in bug report interpretation. This enables LLMs to extract S2R entities more accurately for the target bug report. To overcome the second challenge, BugRepro incorporates app-specific knowledge by exploring the app’s GUI to improve the LLM-guided bug reproduction process further. When LLMs encounter challenges, such as difficulty in determining which UI component to interact with in the reproduction process, BugRepro extracts UI transition graphs(UTG) that model the changes and interactions of the app state on different screens. These graphs are then fed into the LLM to guide its decision-making, addressing the limitation of existing LLMs—their lack of dynamic behavior modeling in app reproduction. This ensures that the reproduction follows the correct sequence of actions and app state transitions.
We conduct extensive experiments on 151 real-world bug reports to evaluate the performance of BugRepro. First, we evaluate the accuracy of S2R extraction and compare BugRepro with two state-of-the-art techniques (i.e., RecDroid and AdbGPT). Experimental results show that BugRepro’s extraction accuracy is 7.57 to 28.89 percentage points higher than prior methods. Then, we evaluate the effectiveness of bug replay. The experimental results show that RecDroid and AdbGPT successfully reproduced 38 and 55 bugs, while BugRepro can reproduce 96 bugs, achieving an improvement of 152.63% and 74.55%, respectively. Furthermore, BugRepro reduces the bug reproduction time to 124.7 seconds on average, compared to 534.9 seconds for RecDroid and 125.6 seconds for AdbGPT. Finally, ablation studies confirm that each component of BugRepro contributes to the overall performance.
@@
  • BugRepro effectively addresses two key challenges in Android bug reproduction: the lack of domain-specific knowledge, the complexity of modern applications with multi-page structures and diverse functionalities.
@@
2 Background and Motivation
2.1 Large Language Models
2.2 Retrieval-Augmented Generation
2.3 Bug Reports and App UI Elements
@@ 2.4 Motivation
Table 1 provides examples of real-world bug reports, and Table 2 presents a comparison of the corresponding S2R entity extraction results. According to the result shown in Table 2, it is obvious that during the S2R entity extraction stage, prior approaches often fail to extract S2R entities when the action verbs are implicit or absent in the sentence. Moreover, these methods frequently fail to extract entities when sentences contain multiple components (e.g., “Secret field” or “other required fields”) or actions (e.g., “enter” or “fill”) of the same type, resulting in incomplete or ambiguous extraction results.
Table 1: Examples of Bug Reports
@@
During the replay stage, prior methods [57, 16] fail to address scenarios requiring multi-page flows or complex functionalities. According to the algorithms used in prior approaches, during the exploration process, if a bug report skips the description of a component on the current screen and directly mentions a component on the next screen, the system tends to perform a “back” operation to return to the previous screen. For example, a bug report says “Select About”, but on the current screen, there is no component labeled “About”. We need to select any component from this screen first, and then we can see the “About” component. When prior methods encounter this situation, they simply choose the “back” action, which never leads to the correct screen. This often leads to incorrect looping behavior, ultimately causing the failure of bug reproduction.
To enhance bug reproduction efficiency, we adopt the RAG approach to overcome issues in the S2R entity extraction stage, and leveraging LLMs with GUI-derived app knowledge for guided replay through UTG construction and interface analysis.
@@ 3 Approach
Our method is structured into two distinct phases and comprises three core components. The first phase is RAG-enhanced S2R entity extraction, we employ RAG to enhance the accuracy of S2R entity extraction (§ 3.1.2). When a new bug report is submitted, it is first segmented into individual sentences. For each sentence, the system automatically retrieves semantically similar sentences and their corresponding labels (which include all associated S2R entities) from the RAG database—an example is shown in Table 3. Then, the original bug report along with the retrieved labeled examples is fed into LLM, which is guided by the examples to analyze and extract the S2R entities corresponding to the new bug report. After extracting the S2R entities, we leverage LLMs to interact with UI components based on these entities and corresponding bug reports to replay the bug. At the same time, based on existing work and our observations, we found that even after applying RAG, challenges remain in extracting components, which may lead to difficulties in correctly matching and operating the corresponding components during the replay process. Therefore, we need to provide the LLM with more UI-related knowledge. The second phase is Exploration-based Replay, which aims to provide app-related knowledge to LLMs, enabling them to gain insights into the app, understand its usage, and make informed decisions. It consists of two primary components: UI exploration based on functionality and LLM guided replay, which continuously interact with each other throughout the process; The functionality-aware UI exploration component is responsible for exploring the UI, and all information will be transmitted to LLM. Note that since UI screens are visual but can not be processed by LLMs directly, we encode the UI screens into text format for LLMs processing following existing work [16]. The LLM guided replay component receives the encoded UI information from the functionality-aware UI exploration module, synthesizes it with inputs from the RAG-enhanced S2R entity extraction module and the original bug report, and then generates guidance for the next interaction step based on this comprehensive context.
After executing each step, Bugrepro checks whether the bug reproduction has succeeded and provides feedback to the LLM. If reproduction fails, it triggers the functionality-aware UI exploration process, allowing the acquisition of additional domain-specific knowledge to enhance the LLM’s decision-making. During the LLM’s decision-making process, when LLM faces difficulties determining the next step (§ 3.2.1), BugRepro augments its decision-making by integrating app-specific knowledge, which encompasses both the UI exploration and the synthesis of app-related information. This process involves analyzing the app’s UI elements and interactions to construct the Synthesized Functionality Table (Table 5), which consists of three components: <Synthesized functionality, UI states, UI elements>. Once equipped with more comprehensive domain-specific knowledge, the LLM is able to make more accurate and effective decisions in response to the current situation.
@@ 3.1.2 RAG-enhanced S2R entity extraction
After defining the available actions and their corresponding primitives, BugRepro incorporates the retrieved examples into the prompt to further assist the LLM. Specifically, BugRepro includes each retrieved sentence along with its corresponding annotated S2R entity, formatted as in the following examples for S2R entity extraction. The sentence is “Click on the add to option and select playlist.”, the extracted S2R entities are: 1. [Tap] [add to option], 2. [Tap] [playlist].
@@ 3.2.1 UTG-Driven Functional Exploration
UI exploration is triggered when the LLM encounters uncertainty regarding which UI component to interact with. Therefore, BugRepro begins the exploration by analyzing the UTG starting from the page where the difficulty arises, denoted as $p$. The UTG, created by the UI Automator, provides essential information about the app, including the relationships between UI states and the presence of various UI elements on each screen. By examining the functionalities of these elements, we can uncover the tasks that can be performed within the app and identify the UI components necessary to execute them. Therefore, BugRepro parses the UI states and elements in the UTG, querying LLMs to extract their respective functions.
@@ 3.2.2 Exploration-enhanced Prompting
However, when the LLM struggles to make the correct decision—such as when it cannot determine which UI component to interact with, or when exploration based on the LLM’s guidance fails to reproduce the bug and the LLM does not update its decision after revisiting the page. In such cases, we provide additional app-related knowledge obtained from the previous exploration to assist the LLM. Specifically, given a problematic page, we supply the LLM with the functionality of all UI components on that problematic page and the functionalities of the new UI screens they can navigate to. This additional information enables the LLM to decide which UI component to interact with and what action to take. At this stage, the prompt includes the following components: (1) the bug report text, (2) the extracted S2R entity, (3) the encoded text of the current UI screen, and (4) the app-related knowledge, which includes the synthesized functionalities (§ 3.2.1) of each UI component within current UI page, and the UI function table, which includes a mapping of functionalities associated with UI pages accessible through specific components on the current page.
@@ 4.2 Implementation and Environment
We implement BugRepro in Python, and we leverage all-MiniLM- L12-v2 model as the retrieval model used in the RAG component, which is based on the Dense Passage Retrieval (DPR) [30] architecture and uses a bi-encoder to encode the queries and documents into dense vectors. We utilize the DeepSeek-V3 model as the underlying LLM in BugRepro, accessed via its API service [13]. To address potential verbosity in the LLM’s output—such as repeated questions or chain-of-thought reasoning—we adopt a filtering mechanism that removes non-JSON content, retaining only valid JSON segments enclosed in square brackets. For interacting with UI widgets on the device, we leverage UI Automator2 [39] as the execution engine. We set the UTG exploration depth to 1 and the time budget to 5 minutes to balance effectiveness and efficiency. All experiments are performed on a workstation running Windows 11 Pro 64-bit, equipped with a 13th Gen Intel(R) Core(TM) i7-13700KF CPU (24 cores, base frequency  3.4GHz) and an NVIDIA GeForce RTX 4090 GPU.
@@ 5.1 RQ1: How accurate is BugRepro in extracting S2R entities?
In the table, BugRepro outperforms both ReCDroid and AdbGPT in all dimensions, with 7.57% to 11.69% higher accuracy than AdbGPT. For instance, BugRepro records 69.49% accuracy in extracting action types, compared to 61.92% by AdbGPT and 45.50% by ReCDroid. Similarly, in identifying target components, BugRepro achieves 33.87%, outperforming AdbGPT (28.77%) and ReCDroid (4.98%). This trend persists across input values and scroll directions, underscoring BugRepro’s consistent advantage.
@@
RQ1 summary: Integrating RAG into our approach clearly demonstrates that BugRepro outperforms baseline methods in S2R extraction, achieving accuracy improvements of 7.57% to 11.69% over AdbGPT (state-of-the-art baselines) across all dimensions. This fully confirms that RAG not only provides contextualized examples, enabling BugRepro to effectively handle ambiguous or complex bug reports, but also enhances contextual understanding, precision in multi-step scenarios, and robustness to linguistic variations.
@@ 5.4 RQ4: How does the robustness of each method vary across different language models?
To answer RQ4, we evaluate BugRepro’s performance using both GPT-4 and our default DeepSeek model to assess its generalization capability. Table 10 presents the results, demonstrating consistent S2R entity extraction performance across both models, with DeepSeek only showing slight disadvantages in Step (87.85% vs. 87.61%) and Component (33.87% vs. 33.76%) extraction. For replay performance, DeepSeek achieves a higher NSR of 96 compared to GPT-4’s 86, though with a longer average reproduction time (124.7s vs. 83.0s).
@@
6 Threats to Validity
The structured-output reliability threat mainly lies in the selection of evaluation metrics and the inherent randomness in LLM outputs. LLMs frequently produce unstructured responses, which impedes their integration with structured code and automation pipelines. To address these challenges, we initially adopt approaches from prior work [17], which attempt to enforce output formatting through prompt examples (e.g., “1. [Tap] [‘bookmark’]”). However, such efforts have not fully standardized the responses. As illustrated in Table 11, LLM outputs often display inconsistencies in format, structure, and verbosity, thereby complicating automation and diminishing accuracy.
To confront this challenge, our approach employs LangChain to enforce output constraints and ensure standardized formatting. The standardized outputs are shown in Table 6. The refinement of the output format further boosts the performance of BugRepro, enforces structured action output an ensure consistency and automation readiness.